OpenStudy (anonymous):

f(x)=5x^2+2 what is the vertex?

OpenStudy (anonymous):

since there is no \(x\) term the first coordinate of the vertex is \(0\)

OpenStudy (anonymous):

@satellite73 would it be 0,2?

OpenStudy (anonymous):

and the axis of symmetry is 2?

OpenStudy (anonymous):

ohh okay thanks! to graph it, do i just plug any numbers into x to get y value?

OpenStudy (anonymous):

just 0?

OpenStudy (anonymous):

@Chad123

OpenStudy (anonymous):

Let me start all over again

OpenStudy (anonymous):

Y=5x^2+2 Point (x,y) is your vertex Let a=5 and b=0 since there is no term for b X=-b/2a = -0/2(5) = 0/10 = 0 X=0 is your line of symmetry To find your vertex replace x with 0 in your function and solve for y Y=5(0)^2+2 Y=2 (0,2) is your vertex

OpenStudy (anonymous):

ohh that makes sense :) thanks so much! and is there a specific formula to find the min/max?